Status of the receiver transformer repair:

Some definitions:
  • PPM crate 0 corresponds to EMB-A side
  • PPM crate 1 corresponds to EMB-C side
  • PPM crate 2 corresponds to EMEC-A side
  • PPM crate 3 corresponds to EMEC-C side
  • PPM crate 4 corresponds to HECFCAL-A side
  • PPM crate 5 corresponds to HECFCAL-C side
  • PPM crate 6 corresponds to TILE-A side
  • PPM crate 7 corresponds to TILE-C side
  • Receiver slots are from 6 to 21
  • Receiver input channels start from 0 (0 to 15) for each cable.
  • Input cable numbering: 1 to 4 from top to bottom.

The problematic channels which were found after combined tests with L1CaloGroup are:

SubDet eta phi Receiver# Channel# comment Repair Status
EMBA 0.2,0.4 5.0,5.2 18 2 & 3 of cables 3 and 4 4 strange weak signal (belongs to one MCM) More likely MCM problem (receiver was checked), status unknown.
EMBA 0.7,0.8 5.1, 5.2 18 7 of cable 4 very weak signal TR. (B5) problem --> fixed
EMBC -1.3,-1.2 1.5 - 1.6 9 12 of cable 4 no signal N/A (Receiver to be checked)
EMBC -1.4,-1.3 1.8 - 1.9 10 13 of cable 4 no signal TR. (A7) problem --> fixed
EMBC -0.4,-0.3 2.0 - 2.1 11 3 of cable 2 very weak signal TR. (D15) problem --> fixed
EMBC -0.1,-0.0 2.0 - 2.1 11 0 of cable 2 very weak signal TR. (D10) problem --> fixed
EMBC -0.1,-0.0 3.3 - 3.4 14 0 of cable 1 weak signal TR. (D9) problem --> fixed
EMBC -1.4,-1.2 2.4 - 2.6 12 2 & 3 of cables 1 and 2 four strange channels (one MCM) spike in saturation More likely MCM problem, status unknown.
TILEA 1.1,1.5 0.2 - 0.3 6 11-14 of cable 3 four channels with known Tile front end problem Tile Problem, status unknown.
TILEA 1.3,1.4 1.2 - 1.3 9 13 of cable 1 no signal Not a receiver problem, status unknown.
TILEA 1.0 ,1.5 1.3 - 1.4 9 10-14 of cable 2 5 channels with no signal Not a receiver problem, status unknown.
TILEA 1.1,1.2 1.4 - 1.5 9 11 of cable 3 no signal Not a receiver problem, status unknown.
TILEA 1.3,1.4 1.4 - 1.5 9 13 of cable 3 no signal Not a receiver problem, status unknown.
TILEA 1.0,1.5 2.1 - 2.2 11 10-14 of cable 2 5 channels with no signal Not a receiver problem, status unknown.
TILEA 1.1,1.2 4.6 - 4.7 17 11 of cable 3 no signal Not a receiver problem, status unknown.
TILEA 1.3,1.4 5.1 - 5.2 18 13 of cable 4 no signal Not a receiver problem, status unknown.
TILEA 1.4,1.5 5.1 - 5.2 18 14 of cable 4 no signal Not a receiver problem, status unknown.
TILEC -1.6,-1.4 3.8 - 4.0 15 14,15 of cables 3,4 four channels with high pedestal (one MCM) More likely MCM problem, status unknown.

The following channels we found dead but were not reported above:

SubDet eta phi Receiver# Channel# comment Repair Status
EMBC -1.3,-1.2 4.6,4.7 17 12 of cable 4 small signal TR. (A6) problem --> fixed

The repair procedure consists in removing the receiver board and make DC tests (passive tests) of all the transformers. And these are channels that had bad transformers although they were not found dead during the active tests.

SubDet eta phi Receiver# Channel# comment Repair Status
EMBC -0.2,-0.1 3.5,3.6 14 1 of cable 3 was not found during active tests TR. (D4) problem --> fixed
EMBC -0.9,-0.8 4.4,4.5 17 8 of cable 2 was not found during active tests TR. (B10) problem --> fixed

The following channels are not used but fixed anyway:

SubDet Receiver# comment Repair Status
TILEA 18 Not used TR. (D6) problem --> fixed

Problems found/fixed on Dec 2007:

SubDet eta phi Receiver# Channel# comment Repair Status
EMBA 0.0,0.1 1.7,1.8 10 0 of cables 2 no signal TR. (A12) problem --> fixed
EMBA 0.1,0.2 2.4,2.5 12 1 of cable 1 half signal TR. (B6) problem --> fixed
EMBA 0.7,0.8 3.4,3.5 14 7 of cable 3 half signal TR. (A14) problem --> fixed
EMBA 0.4,0.5 4.8,4.9 7 4 of cable 3 half signal Not receiver problem (receiver checked)
EMBA 1.1,1.2 1.1,1.2 8 11 of cable 3 half signal Not receiver problem (receiver checked)
EMBA 0.6,0.7 1.5,1.6 9 6 of cable 3 half signal Not receiver problem (receiver checked)
EMBA 0.2,0.3 & 0.5-0.6 1.7,1.8 & 1.8,1.9 10 2 of cable 1 and 5 of cable 3 half signal Not receiver problem (receiver checked)
EMBA 0.2,0.3 & 0.3-0.4 5.0,5.1 & 5.1,5.2 18 2&3 of cable 3 and 2&3 of cable 4 no signal Looks like a problem on the detector side since we could not see signals through the monitoring board.(to be double checked when we get FEC back.)
EMBC -1.2,-1.1 & -1.0,-0.9 & -0.1,0. 4.8-4.9 & 3.2-3.3 6 0&9 of cable 3 and 11 of cable 4 half signal TR. (B5, B2, D4) problems --> fixed
EMBC -0.6,-0.5 5.3-5.4 19 5 of cable 1 half signal TR. (C8) problem --> fixed

-- ChaoukiB - 01 Sep 2007

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r6 - 2007-12-13 - ChaoukiB
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback